solve by substitution method
4m+n=10
m-4n=11
and
6x+7y=-13
-4x+y=37

Answer by checkley77(12844)

4m+n=10 or n=10-4m
m-4n=11 replace n with (10-m) & solve for m.
m-4(10-4m)=11
m-40+16m=11
17m=11+40
17m=51
m=51/17
m=3 ans.
---------
n=10-4*3
n=10-12
n=-2 ans.
------------
6x+7y=-13
-4x+y=37 multiply by -7 & add.
28x-7y=-259
6x+7y=-13
---------------
34x=-272
x=-272/34
x=-8 ans.
6*-8+7y=-13
-48+7y=-13
7y=-13+48
7y=35
y=35/7
y=5 ans.